4(a) For the use of space, a warehouse owner has been offered the following contract: Year 1: $2,000 at the beginning of year 1 plus $2,000 at the end of year 1.
Years 2 to 6: $2,000 per year (end-of-year payments) for years 2 to 6.
Years 7 to 15: $3,000 for year 7 and for the following years, an increase of $1000/year (end of year payments) (i.e. year 7 payment will be $3,000, for year 8 the payment will be $4,000, etc.).

If the warehouse owner can earn 6% per year on the money:
(i) What is the present worth of payments to be received for years 1 to 6? (ii) What is the present worth of payments to be received for years 7 to 15?
Answer #1

Present worth of payments from year 1 to 6 =2000+2000/1.06+2000/1.06^2+...+2000/1.06^6=2000*5.790=10250.46

PW of payments from year 7 to 15

3000/1.06^7+4000/(1.06^8)+(5000/1.06^9)+...+11000/1.06^15

Present worth of payments from year 7 to year 15

PW=3000/(1.06^i =31710.42